yourgrandpa:

jessy took out loan of $16,500 over 5 years at 4.2% interest to purchase a new car. how much will she pay in interest on the loan. A)$693 B)$3,465 C)$13,035 D)$34,650

yourgrandpa:

help needed!!!!

KjSaif:

Well first, do you know the formula for percent? @yourgrandpa

yourgrandpa:

yes

KjSaif:

Ok cool, since we have the percent and the x amount (16,500) we can use this. Y = P% * X

yourgrandpa:

ok

xxDeppressionxx:

i thought that we did 16,500 ÷ 5 4.2 = ? Am i right ?

xxDeppressionxx:

Or am i wrong?

KjSaif:

Ok, the first step would be converting the percent to a decimal.

KjSaif:

What would 4.2 divided by 100 =?

yourgrandpa:

0.042

KjSaif:

Correct. Now that we have it as a decimal we can use that formula I mentioned earlier. Y = P% * X

yourgrandpa:

ok

KjSaif:

It would look like this: 

KjSaif:

uhh that was ugly LOL, but yes, 0.042 x 16500. What does this equal?

yourgrandpa:

A)$693??

yourgrandpa:

thanks (: alot

AZ:

So this is a simple interest calculation

AZ:

The formula that we would have to use is A = P(1 + rt) A would be the total amount paid including the interest P is our starting off amount which is 16,500 in this case r is the interest rate or 4.2% t is time which is 5 years

AZ:

You will find that A is not the correct answer

yourgrandpa:

ok then what is it??? now I'm confused

AZ:

We're dealing with interest and in this case it's about simple interest

yourgrandpa:

ok

AZ:

You have almost all of the numbers, we need to solve for A before we can get your final answer. A = P(1 + rt) We said P = 16500 r = 0.042 (that's 4.2% as a decimal like you calculated earlier) t = 5 So what is A = 16,500 * (1 + (0.042*5)) so first, what is 0.042*5 = ?

yourgrandpa:

0.210

AZ:

Good., now we have A = 16,500 * (1 + 0.21) add 1 to it 1 + 0.21 = ?

yourgrandpa:

1.21

KjSaif:

OH, I see now. What you and I did @yourgrandpa, was only to find the interest value. I forgot to multiply the initial interest times the amount of years. So we can start from where we were at (693) and multiply it by 5. Isn't this correct? @AZ

yourgrandpa:

ohhh ok

AZ:

Well yes, but this will only work for simple interest questions. If it's a compound interest problem, that wouldn't work.

yourgrandpa:

ok

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Good., now we have A = 16,500 * (1 + 0.21) add 1 to it 1 + 0.21 = ? \(\color{#0cbb34}{\text{End of Quote}}\) 1.21

AZ:

multiply that with 16500

yourgrandpa:

19965.00

AZ:

So that's the TOTAL amount you will pay in 5 years with interest. So to calculate how much interest you pay, you have to just subtract the initial amount from it 19,965 - 16,500 = ?

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ So that's the TOTAL amount you will pay in 5 years with interest. So to calculate how much interest you pay, you have to just subtract the initial amount from it 19,965 - 16,500 = ? \(\color{#0cbb34}{\text{End of Quote}}\) 3465

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa 3465 \(\color{#0cbb34}{\text{End of Quote}}\) And that's your final answer!

KjSaif:

Ohh, so now I see how you got your answer @AZ thanks for the heads up earlier.

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa 3465 \(\color{#0cbb34}{\text{End of Quote}}\) And that's your final answer! \(\color{#0cbb34}{\text{End of Quote}}\) thanks(:

AZ:

My pleasure!